# DreamForge: Motion-Aware Autoregressive Video Generation for Multiview Driving Scenes

### v2

@@ 4.2 Qualitative Results
We compare our method with the MagicDrive baseline [17] to assess qualitative results. Fig. 2 shows that our method produces more geometrically accurate multiview images due to perspective guidance. Additionally, Fig. 5 illustrates that our DreamForge performs better in accurate object generation and maintaining consistency in street views, highlighting the effectiveness of the proposed object-wise position encoding. We further explore the effect of ego poses on background changes. As shown in Fig. 6, by modifying the ego car’s pose (from “stop” to “move forward”), we observe that the generated background exhibits significant changes, demonstrating that the network can effectively extract motion cues from the ego poses. Moreover, our method supports additional control conditions, such as road layouts, 3D bounding boxes, and text prompts, enabling more diverse scene generation. Please see the Appendix A.3 for more details.
@@
Long Multiview Videos. By utilizing the motion-aware autoregressive generation pipeline, our method is capable of generating long multiview videos (exceeding 200 frames) with a 7-frame model. Fig. 7 visualizes a long video generated by our 7-frame model. By conditioning on road layouts and 3D bounding boxes, our model can produce videos with over 200 frames at 12 Hz. These results demonstrate that our model is highly 3D-controllable and maintains high video fidelity and consistency across the generated frames. The ability to generate extended video sequences in an autoregressive manner underscores our model’s potential for realistic autonomous driving simulation.
@@ 4.3 Ablation Study
Effectiveness of Object-wise Position Encoding. We further validate the impact of the OPE module. The detailed ablation results are presented in Table 6. When the object position embeddings are directly fed into ControlNet, the FID values decrease by 0.59, while the object mAP and NDS are boosted by 0.63 and 0.34 points, respectively. Further incorporating object position embeddings into the self-attention module of the denoising U-Net brings an improvement of 1.17 mAP and 1.44 NDS. The FID values are also decreased by 0.83. The gains highlight the enhanced foreground generation capability of the proposed OPE module, which introduces only a slight increase in parameters due to the reuse of the self-attention layer. Please refer to the Appendix A.3 for more visual comparisons.
Effectiveness of Motion-aware Temporal Attention. In this section, we discuss the effect of motion-aware temporal attention (MTA). As illustrated in Table 2, by utilizing MTA, our method achieves a lower FVD and improved object mAP and segmentation mIoU, despite generating a 16-frame video using a 7-frame setup, compared to the baseline method trained with 16-frame video clips. We further conducted an ablation study to validate the effectiveness of the local motion module (LMM). The results presented in Table 7 show that LMM can enhance map segmentation by 1.76 points, particularly for the divider and boundary segmentation (with gains of 3.03 and 2.50 points, respectively), while maintaining almost the same object mAP value. This shows local motion implied in feature differences helps the model understand the structure changes.
@@ 5 Conclusion
This paper presents DreamForge, an advanced diffusion-based autoregressive model for long-term 3D-controllable video generation. By utilizing perspective guidance and object-wise position encoding, we enhance the generation quality of street and foreground objects. Additionally, our motion-aware temporal attention effectively captures motion cues, allowing for the generation of long videos (over 200 frames) with a 7-frame model, achieving superior quality compared to the baseline. Finally, we integrate our method with DriveArena to enhance simulation and provide reliable evaluations for vision-based driving agents.
@@ Appendix A Appendix
A.1 Implementation Details
@@
Post-processing strategy Our method does not require a sliding window to maintain coherence between different clips. However, we have experimentally observed that using overlapping frames within a sliding window can slightly enhance generation stability. Additionally, when integrating with DriveArena [13], it is necessary to overlap a few frames to align the output frequency of different components (detailed below). Therefore, we also propose an optional post-processing strategy that utilizes overlapping frames within the sliding windows. Specifically, at the $t$-step of the denoising process for the current video clip, we replace the noised latents $\textbf{Z}^{t}_{T}[\colon N]$ with latents from the previous video clip, denoted as $\sqrt{\overline{\alpha}_{t}}\cdot\widehat{\textbf{Z}}^{0}_{T}[-N\colon]+\sqrt{1-\overline{\alpha}_{t}}\cdot\epsilon_{t}$, before feeding them into the denoising U-Net. Here, $\widehat{\textbf{Z}}^{0}_{T}$ denotes the latents extracted using the VAE encoder, while $\epsilon_{t}$ represents the Gaussian noise at the $t$-step. $\overline{\alpha}_{t}$ is the hype-parameters in the diffusion process. In this way, we ensure that the first $N$ frames of the current video are as consistent as possible with the last $N$ frames of the previous clip for improved coherence.
@@
Training and Inference We train the newly added modules on eight A100 GPUs using the AdamW optimizer [57] with a learning rate of 8e-5. The training process consists of two stages. In the first stage, we train the single-frame version without the motion-aware temporal attention module for 100 epochs with a total batch size of 24. The training objective and hype-parameters are consistent with [17]. In the second stage, we focus solely on training the temporal module for another 100 epochs, using a total batch size of 8. The motion frames are randomly sampled from the previous 5 frames with GT values. For higher-resolution models, we train for 50,000 iterations, initializing with the weights from the smaller-resolution model.
@@
Following the approach outlined in MagicDrive [17], we utilize the UniPC [58] scheduler for 20 steps, applying a CFG of 2.0 to generate the multiview videos. The motion frames are randomly sampled from previously generated video clips. When generating long videos, for the first video clip, we use the single-frame model to generate the initial frame as the motion frames. By default, the length of the overlapping frames in the post-processing strategy is set to 2. Please note that we do not train a new model for different video lengths; all videos of varying lengths are generated using our 7-frame model.
@@
Integration with DriveArena. DriveArena [13] offers a modular platform that can be integrated with any vision-based driving agent for both open-loop and closed-loop simulations. It comprises two key components: (1) Traffic Manager, which processes high-definition maps downloaded from the internet to create diverse urban layouts, manages vehicle movements and traffic flow, and handles collision detection. (2) World Dreamer, a generative model that generates photorealistic multi-view camera images corresponding to the simulation state and adjusts controllable parameters based on specified prompts. All these components exchange data through the network interface. The Traffic Manager runs at 10 Hz while the common vision-based agents such UniAD [7] and VAD [8] take multiview images at 2 Hz. Therefore, it is necessary to synchronize the Traffic Manager, our 7-frame DreamForge, and the driving agents. We utilize a queue of length 7 to cache the data from the Traffic Manager, which is sent to our DreamForge at 2 Hz. In this manner, DreamForge receives 7-frame data each time, with the previous 2 frames overlapping with those from the last iteration. Subsequently, the last frame is taken as the keyframe and sent to the driving agents for planning. In the open-loop mode, the Traffic Manager generates trajectory for ego vehicles. While in closed-loop mode, the trajectory outputted by the driving agents at 2 Hz, consisting of six path points over the next 3 seconds, is interpolated to create a 10 Hz trajectory, which is then directly used for ego vehicle control. Through a motion-aware autoregressive generation paradigm, our DreamForge supports long-term multi-view generation. However, it is inevitable that accumulated errors gradually arise during the iterative process of the simulator. To alleviate the above issues, we empirically used motion frames sampled from the previous clip as conditions for the single-frame version, refining these frames to reduce potential accumulated errors.
A.2 Dataset and Metrics
@@
Dataset. We utilize the nuScenes dataset [1] to train our controllable multiview street view video generation model DreamForge. The nuScenes dataset provides 6 camera views at 12 Hz, offering a 360-degree perspective of the scenes. It includes 750 scenes for training and 150 scenes for validation, encompassing different cities and a variety of lighting and weather conditions, such as daytime, nighttime, sunny, cloudy, and rainy scenarios. Since the nuScenes dataset only provides annotations at 2 Hz, we employ ASAP [59] to generate interpolated annotations at 12 Hz. Additionally, we annotated each scene using GPT-4, providing detailed descriptions that include elements like time, weather, street style, road structure, and appearance. These descriptions serve as conditions for text input.
@@
Metircs. We use FID [60] and FVD [61] to assess the quality of the generated images and videos. Additionally, we evaluate the sim-to-real gap by measuring performance on the generated scenes in downstream tasks, including 3D object detection (mAP and NDS), BEV segmentation (mIoU), and end-to-end planning (L2 and Collation rate). Following [62, 63, 13], we employ the PDM Score (PDMS) and Arena Driving Score (ADS) to evaluate the performance of driving agents in both open-loop and closed-loop modes within the DriveArena simulator [13]. PDMS [62] assesses the trajectory output at each timestep, incorporating penalties for driving without collisions (NC) with road users and compliance with the drivable area (DAC). It also includes a weighted average of factors such as ego progress (EP), time-to-collision (TTC), and comfort (C). Building on PDMS, ADS [13] is calculated by multiplying the modified PDMS by the route completion score [63].
A.3 More Results
Quantitative Results.
@@ A.3 More Results
We also utilize UniAD [7] as an evaluator for the generated scenes to compare various metrics, including 3D object detection, BEV map segmentation, and planning, as illustrated in Table 8. We can see that, compared to the state-of-the-art method MagicDrive, our method outperforms it in nearly all metrics, except for a slight lag in the collision rate. Interestingly, we found that when we increase the input resolution, the collision rate in the scenes generated by our method is lower than that observed in the real data. Additionally, the L2 error is also smaller when using the same input resolution of $224\times 400$.
@@
Figure 12: Visualizations featuring various text prompts for different weather conditions, such as sunny, rainy, and nighttime. For a better view, we visualize several sampled keyframes from the generated videos.
Visualizations. We provide more cases in Fig. 8 to demonstrate that our DreamForge performs better in accurate object generation and maintaining consistency in street views. The visualizations illustrate that our method produces buses, trucks, and people with improved appearance, particularly in the cross-view area. We also offer visualizations that utilize the road layouts and 3D bounding boxes generated by DriveArena [13]. The results, presented in Fig. 9 and Fig. 10, demonstrate that our method can generate controllable urban scenes by modifying the road layouts, as well as the bounding boxes of objects. Furthermore, Fig. 12 demonstrates how DreamForge alters scenes to replicate different weather conditions and times of day, emphasizing its controllability in modifying scene descriptions. Finally, we illustrate the visualizations of the simulations within DriveArena. Our DreamForge receives sequential data from the Traffic Manager in DriveArena and sends generated scenes of keyframes to the driving agents at 2 Hz, as stated in the implementation details. Without loss of generality, we present cases where the driving agents UniAD [7] and VAD [8] operate in an open-loop mode in Fig. 11. The results indicate that predictions from the driving agents regarding the road network and vehicle tracking are fundamentally accurate, with UniAD demonstrating superior perception results on the generated scenes.

### v3

@@ 4.2 Qualitative Results
We compare our method with the MagicDrive baseline [17] to assess qualitative results. Fig. 2 shows that our method produces more geometrically accurate multiview images due to perspective guidance. Additionally, Fig. 5 illustrates that our DreamForge performs better in accurate object generation and maintaining consistency in street views, highlighting the effectiveness of the proposed object-wise position encoding. In addition, our method supports additional control conditions, such as road layouts, 3D boxes, and text prompts for generating diverse scenes, including smooth weather transitions. We also observe that ego pose influences the generated background, demonstrating the effectiveness of motion cues from ego poses. Please see Appendix A.4.2 for more details.
@@
Long Multiview Videos. Our motion-aware autoregressive generation pipeline enables the synthesis of long multiview videos (over 200 frames) using a model trained on short sequences. As shown in Fig. 6, our model, conditioned on road layouts and 3D bounding boxes, produces videos at 12 Hz with high 3D controllability, fidelity, and frame consistency. This capability highlights its potential for realistic autonomous driving simulations. Please see Appendix A.4.2 and the project page for more long video demo.
@@ 4.3 Ablation Study
Effectiveness of Different Base Model. To demonstrate that our proposed method can adapt to different base models, we conduct a detailed ablation study, as shown in Table 3. We examine two configurations of the base model: SD V1.5 [56] and DiT [28, 66] with the 3D VAE [57]. The results indicate that our model consistently improves both mAP and mIoU across both configurations. Furthermore, using autoregressive generation achieves comparable FVD to the baseline when employing the same base model. Notably, employing DiT with the 3D VAE significantly enhances FVD, demonstrating the effectiveness of the temporal connections in the 3D VAE and the stronger temporal modeling capabilities of DiT. However, we also observe that using DiT with the 3D VAE results in a decline in mAP performance for object detection. We attribute this to the temporal downsampling applied to bounding boxes, which may introduce positional ambiguity, especially for dynamic objects, thereby affecting fine-grained controllability. This will be a focus for future exploration.
@@
Effectiveness of Object-wise Position Encoding. We further validate the impact of the OPE module. The detailed ablation results are presented in Table 7. When the object position embeddings are directly fed into ControlNet, the FID values decrease by 0.59, while the object mAP and NDS are boosted by 0.63 and 0.34 points, respectively. Further incorporating object position embeddings into the self-attention module of the denoising blocks brings an improvement of 1.17 mAP and 1.44 NDS. The FID values are also decreased by 0.83. The gains highlight the enhanced foreground generation capability of the proposed OPE module, which introduces only a slight increase in parameters due to the reuse of the self-attention layer. Please refer to Appendix A.4.2 for more visual comparisons.
Effectiveness of Motion-aware Temporal Attention. In this section, we analyze the impact of motion-aware temporal attention (MTA). As shown in Table 2, our method, leveraging MTA, achieves lower FVD and higher object mAP and segmentation mIoU compared to the baseline trained on 16-frame clips, even when generating 16-frame videos from a 7-frame setup. This highlights MTA’s effectiveness in capturing temporal dynamics. Additionally, an ablation study on the local motion module (LMM) in Table 8 reveals its ability to improve map segmentation by 1.76 points, particularly for the divider and boundary segmentation (gains of 3.03 and 2.50 points, respectively), while preserving object mAP. These results suggest that LMM effectively leverages local motion to enhance structural understanding.
@@ 5 Conclusion
This paper introduces DreamForge, an advanced diffusion-based autoregressive model for long-term 3D-controllable video generation. By incorporating perspective guidance and object-wise position encoding, we enhance the quality of street and foreground object generation. Additionally, our motion-aware temporal attention effectively captures motion cues, enabling the generation of long videos (over 200 frames) with a model trained on short sequences, outperforming the baseline in quality. Finally, we integrate our method with DriveArena to improve simulation and provide reliable evaluations for vision-based driving agents.
@@ Appendix A Appendix
A.1 More Related Works
A.1.1 Diffusion-based conditional generation
Diffusion models have revolutionized generative tasks in both the image and video domains [67, 68, 69, 28, 70]. In the realm of image generation, models like Stable Diffusion [56], PixArt [66], and Flux [71] are capable of generating high-quality images, whereas video diffusion techniques, exemplified by Stable Video Diffusion [27], OpenSora [28] and Cogvideox [57], mitigate challenges associated with temporal consistency and motion dynamics. While traditional Text-to-Image (T2I) and Text-to-Video (T2V) methods [72, 56, 66, 73, 74, 57, 28] often struggle to provide precise control over the generated content, ControlNet [52] merges and addresses this limitation by training a control network that copies parts of the pre-trained main model to introduce control conditions, such as edge maps, segmentation masks, and poses.
In the field of autonomous driving, precise control over video generation plays a vital role in developing realistic simulations. Recently, diffusion-based controllable methods such as MagicDrive [17], DrivingDiffusion[22], and Panacea [21] have emerged for street-view scene generation. These approaches integrate 3D bounding boxes, Bird’s Eye View (BEV) maps, ego trajectories, and camera poses to synthesize multi-view street scenes. To take advantage of the strong spatiotemporal modeling of transformers [66, 28], MagicDriveDiT [64] integrate the DiT architecture [75, 47] with 3D VAEs [28, 57] to manage spatiotemporal latent representations. Different from the above methods, we propose a motion-aware autoregressive architecture, which introduce perspective guidance and object-wise position encoding to improve controllability and motion-aware temporal attention to improve temporal coherence and seamless video generation. It can well adapt to various generative base models, highlighting its broad applicability to the autonomous driving community.
A.2 Implementation Details
A.2.1 Post-processing strategy
@@
We have experimentally observed that using overlapping frames within a sliding window can slightly enhance generation stability. Furthermore, integrating with DriveArena [13], it is necessary to overlap a few frames to align the output frequency of different components (detailed below). Therefore, we also propose an optional post-processing strategy that utilizes overlapping frames within the sliding windows. Specifically, at the $t$-step of the denoising process for the current video clip, we replace the noised latents $\textbf{Z}^{t}_{T}[\colon N]$ with latents from the previous video clip, denoted as $\sqrt{\overline{\alpha}_{t}}\cdot\widehat{\textbf{Z}}^{0}_{T}[-N\colon]+\sqrt{1-\overline{\alpha}_{t}}\cdot\epsilon_{t}$, before feeding them into the denoising U-Net. Here, $\widehat{\textbf{Z}}^{0}_{T}$ denotes the latents extracted using the VAE encoder, while $\epsilon_{t}$ represents the Gaussian noise at the $t$-step. $\overline{\alpha}_{t}$ is the hype-parameters in the diffusion process. In this way, we ensure that the first $N$ frames of the current video are as consistent as possible with the last $N$ frames of the previous clip for improved coherence.
A.2.2 Training and Inference
@@
We train the newly added modules on eight A100 GPUs using the AdamW optimizer [76] with a learning rate of 8e-5. The training process consists of two stages. In the first stage, we train the single-frame version without the motion-aware temporal attention module for 100 epochs with a total batch size of 24. The training objective and hype-parameters are consistent with [17]. In the second stage, we focus solely on training the temporal module for another 100 epochs, using a total batch size of 8. The motion frames are randomly sampled from the previous 5 frames with GT values. For higher-resolution models, we train for 50,000 iterations, initializing with the weights from the smaller-resolution model. For the variants employing DiT [28] and 3D VAE [57] as the base models, we trained from scratch for 300K iterations at a resolution of $224\times 400$, utilizing 8 A100 GPUs with a batch size of 4 per GPU. Subsequently, we proceeded to train at a larger resolution of $448\times 800$ for an additional 50K iterations. Finally, we leveraged the pretrained weights to train the video version at $448\times 800$ resolution for another 50K iterations.
@@
Following the approach outlined in MagicDrive [17, 64], we employ the UniPC [77] scheduler for 20 steps with the base model of SD V1.5 and the rectified flow [78] scheduler for 30 steps with the base model of DiT, applying a CFG (Classifier-Free Guidance) scale of 2.0 to generate the multiview videos. The motion frames are sampled from previously generated video clips. When generating long videos, for the first video clip, we use the single-frame model to generate the initial frame as the motion frames. By default, the length of the overlapping frames in the post-processing strategy is set to 2. Note that we do not train a new model for different video lengths; all videos of varying lengths are generated using the model trained in short sequences.
A.2.3 Integration with DriveArena.
@@
DriveArena [13] offers a modular platform that can be integrated with any vision-based driving agent for both open-loop and closed-loop simulations. It comprises two key components: (1) Traffic Manager, which processes high-definition maps downloaded from the internet to create diverse urban layouts, manages vehicle movements and traffic flow, and handles collision detection. (2) World Dreamer, a generative model that generates photorealistic multi-view camera images corresponding to the simulation state and adjusts controllable parameters based on specified prompts. All these components exchange data through the network interface. The Traffic Manager runs at 10 Hz while the common vision-based agents such UniAD [7] and VAD [8] take multiview images at 2 Hz. Therefore, it is necessary to synchronize the Traffic Manager, our 7-frame DreamForge, and the driving agents. We utilize a queue of length 7 to cache the data from the Traffic Manager, which is sent to our DreamForge at 2 Hz. In this manner, DreamForge receives 7-frame data each time, with the previous 2 frames overlapping with those from the last iteration. Subsequently, the last frame is taken as the keyframe and sent to the driving agents for planning. In the open-loop mode, the Traffic Manager generates trajectory for ego vehicles. While in closed-loop mode, the trajectory outputted by the driving agents at 2 Hz, consisting of six path points over the next 3 seconds, is interpolated to create a 10 Hz trajectory, which is then directly used for ego vehicle control. Through a motion-aware autoregressive generation paradigm, our DreamForge supports long-term multi-view generation. However, it is inevitable that accumulated errors gradually arise during the iterative process of the simulator. To alleviate the above issues, we empirically used motion frames sampled from the previous clip as conditions for the single-frame version, refining these frames to reduce potential accumulated errors.
Figure 8: By combining autoregressive generation and text prompts for different weather conditions, our model can produce videos that showcase a seamless weather transition within continuous clips. We also find the model still has some limitations such as the accurate transition of light reflection as shown in the red circle, which will be a focus of our future improvement.
A.3 Dataset and Metrics
A.3.1 Dataset.
@@ A.3 Dataset and Metrics
We utilize the nuScenes dataset [1] to train our controllable multiview street view video generation model DreamForge. The nuScenes dataset provides 6 camera views at 12 Hz, offering a 360-degree perspective of the scenes. It includes 750 scenes for training and 150 scenes for validation, encompassing different cities and a variety of lighting and weather conditions, such as daytime, nighttime, sunny, cloudy, and rainy scenarios. Since the nuScenes dataset only provides annotations at 2 Hz, we employ ASAP [79] to generate interpolated annotations at 12 Hz. Additionally, we annotated each scene using GPT-4, providing detailed descriptions that include elements like time, weather, street style, road structure, and appearance. These descriptions serve as conditions for text input.
A.3.2 Metircs.
@@
We use FID [80] and FVD [81] to assess the quality of the generated images and videos. Additionally, we evaluate the sim-to-real gap by measuring performance on the generated scenes in downstream tasks, including 3D object detection (mAP and NDS), BEV segmentation (mIoU), and end-to-end planning (L2 and Collation rate). Following [82, 83, 13], we employ the PDM Score (PDMS) and Arena Driving Score (ADS) to evaluate the performance of driving agents in both open-loop and closed-loop modes within the DriveArena simulator [13]. PDMS [82] assesses the trajectory output at each timestep, incorporating penalties for driving without collisions (NC) with road users and compliance with the drivable area (DAC). It also includes a weighted average of factors such as ego progress (EP), time-to-collision (TTC), and comfort (C). Building on PDMS, ADS [13] is calculated by multiplying the modified PDMS by the route completion score [83].
A.4 More Results
A.4.1 Quantitative Results.
@@ A.4 More Results
We also utilize UniAD [7] as an evaluator for the generated scenes to compare various metrics, including 3D object detection, BEV map segmentation, and planning, as illustrated in Table 9. We can see that, compared to the state-of-the-art method MagicDrive, our method outperforms it in nearly all metrics, except for a slight lag in the collision rate. Interestingly, we found that when we increase the input resolution, the collision rate in the scenes generated by our method is lower than that observed in the real data. Furthermore, the L2 error is also smaller when using the same input resolution of $224\times 400$.
A.4.2 Visualizations.
3D controllability. We provide more cases in Fig. 9 to demonstrate that our DreamForge performs better in accurate object generation and maintaining consistency in street views. The visualizations illustrate that our method produces buses, trucks, and people with improved appearance, particularly in the cross-view area. We also offer visualizations that utilize the road layouts and 3D bounding boxes generated by DriveArena [13]. The results, presented in Fig. 10 and Fig. 11, demonstrate that our method can generate controllable urban scenes by modifying the layout of the roads and the box boundaries of objects. We further explore the effect of ego poses on background changes. As shown in Fig. 7, by modifying the ego car’s pose (from “stop” to “move forward”), we observe that the generated background exhibits significant changes, demonstrating that the network can effectively extract motion cues from the ego poses.
Weather alteration. The examples in Fig. 13 demonstrate how DreamForge transforms scenes to replicate diverse weather conditions and times of day, highlighting its controllability in modifying scene descriptions. Additionally, Fig. 8 showcases a seamless weather transition within continuous video clips. These cases clearly illustrate that the proposed motion-aware autoregressive generation method provides a highly flexible approach to video appearance control. It not only supports the generation of varying weather conditions for the same scene but also facilitates smooth and natural appearance transitions within continuous video sequences. We also find that the model has some limitations, such as its inability to effectively handle changes in light reflection. For instance, as shown in the red circle in Fig. 8 it fails to accurately represent the transition of light reflection on lane lines from nighttime to daytime. Addressing this issue will be our future work.
Long video generation. We present additional case studies in Fig. 14 and Fig. 15 that highlight the 3D controllability of our model, particularly its ability to maintain high video fidelity and temporal consistency across generated frames. These examples illustrate that our model can produce video sequences that are both multiview and temporally coherent, generating realistic content in an autoregressive manner. This capability demonstrates its potential applications in realistic autonomous driving simulation, where accurate and consistent video generation is crucial for the training and testing of autonomous systems.
Simulation within DriveArena. We illustrate the visualizations of the simulations within DriveArena. Our DreamForge receives sequential data from the Traffic Manager in DriveArena and sends generated scenes of keyframes to the driving agents at 2 Hz, as stated in the implementation details. Without loss of generality, we present cases where the driving agents UniAD [7] and VAD [8] operate in an open-loop mode in Fig. 12. The results indicate that predictions from the driving agents regarding the road network and vehicle tracking are fundamentally accurate, with UniAD demonstrating superior perception results on the generated scenes.
@@
Figure 13: Visualizations featuring various text prompts for different weather conditions, such as sunny, rainy, and night. For a better view, we visualize several sampled keyframes from the generated videos.
Figure 14: The visualizations illustrate that our model can produce video sequences that are both multiview and temporally coherent, generating realistic content in an autoregressive manner.
Figure 15: The visualizations illustrate that our model can produce video sequences that are both multiview and temporally coherent, generating realistic content in an autoregressive manner.